# Generative Language Models on Nucleotide Sequences of Human Genes

### v2

@@ 1 Introduction
Generative models have proven their success in different domains, most obviously with ChatGPT [ChatGPT]. Theoretically, they can carry out almost all of the tasks done by a discriminative model. Especially combining different domains such as natural language and programming yielded incredible outcomes like OpenAI Codex [chen2021evaluating], which can write code by only getting natural language as input. DALL-E 2 [ramesh2022hierarchical] is another nice example from a different field. One of its most exciting achievements is replacing an object in an image with another thing without changing the background, called in-painting. Similar work can be done for the combination of any domain and natural language. Although the current version of our study only focuses on genes, the same idea could be applied to whole DNA sequences, and a similar combination with natural language could provide huge benefits. To give concrete examples, we might show a mutated DNA sequence causing disease and prompt it to convert this DNA sequence into the healthy version. Another option could be giving a sequence carrying a specific property like being resistant as input and wanting the program to describe how to integrate that feature into another organism’s DNA so that it has the same property as well. In fact, theoretical limits for what can be done may be very close to the limits of what genomics could do at all. When all these possibilities are taken into account, we had better think carefully about ethical considerations too.
@@ 2 Related Work
Biology benefits from these machines with great potential. Various biological building blocks could be studied much more robustly and quickly with these solid computing machines [yang2020review, li2021bioseq]. Such efforts are sometimes referred to as bioinformatics or computational biology.
@@
Proteins are one of the most fundamental building blocks when it comes to biology. Both discriminative and generative models are used for these critical elements. While McDowall and Hunter [mcdowall2011interpro] provide a tool for automatic classification of proteins using models such as hidden Markov models, Madani et al. [madani2020progen] attempt to develop a generative model for proteins based on big data that can be used for various application domains like medicine.
@@
Since proteins can have different 3D structures, working with them requires special care. In addition, there are studies that focus on the interactions between proteins and various chemicals, as these are of great importance for topics such as drug development. For example, Nagamine and Sakakibara [nagamine2007statistical] use Support Vector Machines, a conventional machine learning algorithm, to predict the probability of interaction for a given protein and chemical pair. On the other hand, Cobanoglu et al. [cobanoglu2013predicting] use probabilistic matrix factorization, which is an alternative method for a closely related task.
@@
DNA and RNA sequences have a massive importance in our lives; consequently, applications related to them are prevalent. The area is not new, even though various facilities in the past were limited. For example, Wang et al. [wang1999new] devised two algorithms, one unsupervised, for classifying DNA. Nonetheless, Nguyen et al. [nguyen2016dna] prefer using convolutional neural networks for classifying DNA even though convolutional neural networks are not commonly used for classifying sequence-like inputs. Furthermore, Qi et al. [qi2020clustering] use an unsupervised clustering method on RNA sequences in addition to supervised applications.
@@
Some studies used classical artificial intelligence methods to solve bioinformatics problems without using machine learning. Altschul et al. [altschul1990basic] use such an approach to measure the similarity of two sequences, as this is a crucial problem for tasks like search. Burge and Karlin [burge1997prediction] use a probabilistic approach to solve tasks such as gene identification. Lowe and Eddy [lowe1997trnascan] use the same technique for RNA-related work with a very low error rate. Also, Trapnell et al. [trapnell2013differential] develop a robust method for gene expression by taking advantage of RNA sequences. Despite the existence of such studies, the integration of machine learning techniques in one form or another is extremely popular nowadays.
@@
Traditional machine learning methods are a serious option for tasks with limited data, as deep learning approaches, their main competitors, need more data to perform better. Rannala and Yang [rannala2003bayes] pursue a statistical method for estimating different species using DNA. Choi et al. [choi2017evaluation] investigate the success of the classical logistic regression algorithm in RNA sequence studies.
@@
As in other fields, deep learning is inevitably very common in bioinformatics because it is very successful when enough data is available for a given task. Lan et al. [lan2018survey] mention several use cases for this, which is expected since deep learning is extremely popular in this field. As an interesting example, Zeng et al. [zeng2023evaluating] deep learning to make predictions about RNA and protein pairs to better understand these structures.
@@
There are some other works using machine learning on proteins. AIPs-SnTCN [raza2023aips] uses a combination of embedding techniques to make classification on peptides. Moreover, cACP-DeepGram [akbar2022cacp] uses combinations of deep learning methods for a cancer-related study.
@@
Lastly, using language models is also a popular topic. For instance, DNABert [ji2021dnabert] uses an approach similar to the classical BERT, where data comes from human DNA for training and various DNA sources for testing instead of natural language in order to solve various tasks related to genomics. While DNABert worked on datasets where samples were DNA sequences, our study was on a smaller scale since our samples were nucleotide sequences of genes.
@@
In addition, Luo et al. [luo2022biogpt] try to do a similar study for the generative part of the coin, yet for biomedical text instead of directly using biological units. However, a successful study which trained a gigantic generative transformer model on structures like DNA or proteins is not available, to the best of our knowledge.
@@
Like natural language processing, biomedical domains try to use language models as much as possible due to their vast potential and proven success in several applications. This circumstance is also called biological language models. It is possible to find very different use cases, ranging from discriminative to generative when the subject is examined in depth [osmanbeyoglu2011n, wang2019high, li2021bioseq, benegas2022dna, gankin2023species, liang2012segmenting].
@@
To provide a summary, Transformer model [vaswani2017attention] provided huge improvements in Natural Language Processing. This is valid for both discriminative models like BERT [devlin2018bert], RoBERTa [liu2019roberta] and generative models like GPT [radford2018improving], GPT-2 [radford2019language], GPT-3 [brown2020language]. Owing to similarity between natural languages and biological languages, they were tried on biological data like DNA and Protein sequences as well, yet it was done mainly for discriminative cases. DNABERT [ji2021dnabert], ProteinBERT [brandes2022proteinbert] and DNABERT-2 [zhou2023dnabert] gave successful results in these domains. These crucial milestones from Natural Language Processing or Bioinformatics are listed chronologically in Table 1.
@@
Transformer [vaswani2017attention] 2017
GPT [radford2018improving] 2018
BERT [devlin2018bert] 2018
GPT-2 [radford2019language] 2019
RoBERTa [liu2019roberta] 2019
GPT-3 [brown2020language] 2020
DNABERT [ji2021dnabert] 2021
ProteinBERT [brandes2022proteinbert] 2022
DNABERT-2 [zhou2023dnabert] 2023
@@ 3 Dataset
The aim of preparing the dataset was to use the nucleotide sequences for genes existing in human beings. The National Center for Biotechnology Information (NCBI) [ncbi], a government institution, provides different databases such as Genome, Gene, and Nucleotide. From the Gene Database, we filtered genes such that only the ones existing in human beings remained. Then we got information like the commonly used name, id used in the database, and nucleotide sequence for all these genes. Nucleotide sequences for a specific gene can be accessed using the Nucleotide database of NCBI, and the dataset we obtained is online [datasets]. Some general statistics about the dataset are given in Table 2 to provide a general overview of what kind of data we worked on.
@@
Next, we had to filter out genes whose sequence length exceeded a certain threshold, since transformer-based models assume a certain maximum length. This threshold was set at 1000, since larger values led to problems in fitting the models to the memory of the computers. Then we used 80% of it for training and 20% of the dataset for testing purposes. The split was done randomly. Some general statistics on the datasets can be found in Table 3. The datasets are available online [datasets].
@@ 4.1 N-gram Language Models
For N-gram language models [shannon1948mathematical], we experimented with six different values of N, starting from one and going up to six. The perplexity values obtained are given in Table 4. Moreover, Natural Language Toolkit (NLTK) library [nltk] was helpful for both using N-grams and calculating perplexity.
@@ 4.1.1 N-gram Language Models with Laplace Smoothing
Then Laplace smoothing [lidstone1920note] is used to avoid this problem. Different N values are tried starting from six and going up to eight. The corresponding perplexity values obtained are given in Table 5.
@@ 4.2 Long Short-Term Memory Based Language Model
After experimenting with N-gram language models, we continued with deep learning-based approaches. The first option was using Long Short-Term Memory [hochreiter1997long] based models. For getting input, a layer for learning embedding and position information was used at the beginning. Then a single or multiple Long Short-Term Memory layers are used depending on hyperparameter configuration. Lastly, the aim was predicting the next word; because of this, a linear layer with the size of vocabulary many units are used with Softmax to get predictions. The model architecture is visualized in Figure 3. Adam was chosen as the optimizer to use.
@@ 4.3 Transformer Based Language Model
Lastly, Tensorflow library [tensorflow], its high-level API Keras [keras] and its NLP tool KerasNLP [kerasnlp] were very beneficial for implementing stuff related to Deep Learning. The best model we obtained had 1,844,988 parameters.
@@
The results obtained for different models based on test set perplexity is visualized in Figure 5 using Matplotlib library [Hunter:2007].
@@ 4.4 Performances On Real Life Tasks
Evaluating the success of the methods is not trivial, so a real-life task could be very useful. For this reason, we tried to find a real-life problem where modeling nucleotide sequences of human genes could be helpful. Distinguishing between real nucleotide sequences of a real gene and a mutated version seemed like an exciting task, since the ability to do such a task well requires some level of understanding of the data we were working on. In fact, what we did is similar to the evaluation of GPT-2 [radford2019language] on tasks like Winograd Schema Challenge [levesque2012winograd], in the sense that our model is trained as a generative model, but it could be adopted to do discriminative tasks when the correct setting is prepared.
@@ 4.4.2 Real Mutation Dataset
Human Gene Mutation Database [stenson2003human, stenson2014human, stenson2020human] was a nice resource with almost no alternative for our goal. Even though the number of mutations we could find for genes in our test set was only 10, which is very limited, it still helped to show that our models could make nice predictions for real examples as well. Since the database is not publicly available, we will, unfortunately, not be able to share the obtained mutation dataset.

### v3

# Musa Nuri İhtiyar, Arzucan Özgür

@@ 1 Introduction
Generative models have proven their success in different domains, most obviously with ChatGPT [32]. Theoretically, they can carry out almost all of the tasks done by a discriminative model. Especially combining different domains such as natural language and programming yielded incredible outcomes like OpenAI Codex [7], which can write code by only getting natural language as input. DALL-E 2 [37] is another nice example from a different field. One of its most exciting achievements is replacing an object in an image with another thing without changing the background, called in-painting. Similar work can be done for the combination of any domain and natural language. Although the current version of our study only focuses on genes, the same idea could be applied to whole DNA sequences, and a similar combination with natural language could provide huge benefits. To give concrete examples, we might show a mutated DNA sequence causing disease and prompt it to convert this DNA sequence into the healthy version. Another option could be giving a sequence carrying a specific property like being resistant as input and wanting the program to describe how to integrate that feature into another organism’s DNA so that it has the same property as well. In fact, theoretical limits for what can be done may be very close to the limits of what genomics could do at all. When all these possibilities are taken into account, we had better think carefully about ethical considerations too.
@@ 2 Related Work
Biology benefits from these machines with great potential. Various biological building blocks could be studied much more robustly and quickly with these solid computing machines [49, 22]. Such efforts are sometimes referred to as bioinformatics or computational biology.
@@
Proteins are one of the most fundamental building blocks when it comes to biology. Both discriminative and generative models are used for these critical elements. While McDowall and Hunter [29] provide a tool for automatic classification of proteins using models such as hidden Markov models, Madani et al. [28] attempt to develop a generative model for proteins based on big data that can be used for various application domains like medicine.
@@
Since proteins can have different 3D structures, working with them requires special care. In addition, there are studies that focus on the interactions between proteins and various chemicals, as these are of great importance for topics such as drug development. For example, Nagamine and Sakakibara [30] use Support Vector Machines, a conventional machine learning algorithm, to predict the probability of interaction for a given protein and chemical pair. On the other hand, Cobanoglu et al. [9] use probabilistic matrix factorization, which is an alternative method for a closely related task.
@@
DNA and RNA sequences have a massive importance in our lives; consequently, applications related to them are prevalent. The area is not new, even though various facilities in the past were limited. For example, Wang et al. [47] devised two algorithms, one unsupervised, for classifying DNA. Nonetheless, Nguyen et al. [31] prefer using convolutional neural networks for classifying DNA even though convolutional neural networks are not commonly used for classifying sequence-like inputs. Furthermore, Qi et al. [34] use an unsupervised clustering method on RNA sequences in addition to supervised applications.
@@
Some studies used classical artificial intelligence methods to solve bioinformatics problems without using machine learning. Altschul et al. [2] use such an approach to measure the similarity of two sequences, as this is a crucial problem for tasks like search. Burge and Karlin [6] use a probabilistic approach to solve tasks such as gene identification. Lowe and Eddy [26] use the same technique for RNA-related work with a very low error rate. Also, Trapnell et al. [45] develop a robust method for gene expression by taking advantage of RNA sequences. Despite the existence of such studies, the integration of machine learning techniques in one form or another is extremely popular nowadays.
@@
Traditional machine learning methods are a serious option for tasks with limited data, as deep learning approaches, their main competitors, need more data to perform better. Rannala and Yang [38] pursue a statistical method for estimating different species using DNA. Choi et al. [8] investigate the success of the classical logistic regression algorithm in RNA sequence studies.
@@
As in other fields, deep learning is inevitably very common in bioinformatics because it is very successful when enough data is available for a given task. Lan et al. [20] mention several use cases for this, which is expected since deep learning is extremely popular in this field. As an interesting example, Zeng et al. [50] deep learning to make predictions about RNA and protein pairs to better understand these structures.
@@
There are some other works using machine learning on proteins. AIPs-SnTCN [39] uses a combination of embedding techniques to make classification on peptides. Moreover, cACP-DeepGram [1] uses combinations of deep learning methods for a cancer-related study.
@@
Lastly, using language models is also a popular topic. For instance, DNABert [19] uses an approach similar to the classical BERT, where data comes from human DNA for training and various DNA sources for testing instead of natural language in order to solve various tasks related to genomics. While DNABert worked on datasets where samples were DNA sequences, our study was on a smaller scale since our samples were nucleotide sequences of genes.
@@
In addition, Luo et al. [27] try to do a similar study for the generative part of the coin, yet for biomedical text instead of directly using biological units. However, a successful study which trained a gigantic generative transformer model on structures like DNA or proteins is not available, to the best of our knowledge.
@@
Like natural language processing, biomedical domains try to use language models as much as possible due to their vast potential and proven success in several applications. This circumstance is also called biological language models. It is possible to find very different use cases, ranging from discriminative to generative when the subject is examined in depth [33, 48, 22, 3, 12, 23].
@@
To provide a summary, Transformer model [46] provided huge improvements in Natural Language Processing. This is valid for both discriminative models like BERT [10], RoBERTa [25] and generative models like GPT [35], GPT-2 [36], GPT-3 [5]. Owing to similarity between natural languages and biological languages, they were tried on biological data like DNA and Protein sequences as well, yet it was done mainly for discriminative cases. DNABERT [19], ProteinBERT [4] and DNABERT-2 [51] gave successful results in these domains. These crucial milestones from Natural Language Processing or Bioinformatics are listed chronologically in Table 1.
@@
Transformer [46] 2017
GPT [35] 2018
BERT [10] 2018
GPT-2 [36] 2019
RoBERTa [25] 2019
GPT-3 [5] 2020
DNABERT [19] 2021
ProteinBERT [4] 2022
DNABERT-2 [51] 2023
@@ 3 Dataset
The aim of preparing the dataset was to use the nucleotide sequences for genes existing in human beings. The National Center for Biotechnology Information (NCBI) [11], a government institution, provides different databases such as Genome, Gene, and Nucleotide. From the Gene Database, we filtered genes such that only the ones existing in human beings remained. Then we got information like the commonly used name, id used in the database, and nucleotide sequence for all these genes. Nucleotide sequences for a specific gene can be accessed using the Nucleotide database of NCBI, and the dataset we obtained is online [18]. Some general statistics about the dataset are given in Table 2 to provide a general overview of what kind of data we worked on.
@@
Next, we had to filter out genes whose sequence length exceeded a certain threshold, since transformer-based models assume a certain maximum length. This threshold was set at 1000, since larger values led to problems in fitting the models to the memory of the computers. Then we used 80% of it for training and 20% of the dataset for testing purposes. The split was done randomly. Some general statistics on the datasets can be found in Table 3. The datasets are available online [18].
@@ 4.1 N-gram Language Models
For N-gram language models [40], we experimented with six different values of N, starting from one and going up to six. The perplexity values obtained are given in Table 4. Moreover, Natural Language Toolkit (NLTK) library [44] was helpful for both using N-grams and calculating perplexity.
@@ 4.1.1 N-gram Language Models with Laplace Smoothing
Then Laplace smoothing [24] is used to avoid this problem. Different N values are tried starting from six and going up to eight. The corresponding perplexity values obtained are given in Table 5.
@@ 4.2 Long Short-Term Memory Based Language Model
After experimenting with N-gram language models, we continued with deep learning-based approaches. The first option was using Long Short-Term Memory [16] based models. For getting input, a layer for learning embedding and position information was used at the beginning. Then a single or multiple Long Short-Term Memory layers are used depending on hyperparameter configuration. Lastly, the aim was predicting the next word; because of this, a linear layer with the size of vocabulary many units are used with Softmax to get predictions. The model architecture is visualized in Figure 3. Adam was chosen as the optimizer to use.
@@ 4.3 Transformer Based Language Model
Lastly, Tensorflow library [15], its high-level API Keras [14] and its NLP tool KerasNLP [13] were very beneficial for implementing stuff related to Deep Learning. The best model we obtained had 1,844,988 parameters.
@@
The results obtained for different models based on test set perplexity is visualized in Figure 5 using Matplotlib library [17].
@@ 4.4 Performances On Real Life Tasks
Evaluating the success of the methods is not trivial, so a real-life task could be very useful. For this reason, we tried to find a real-life problem where modeling nucleotide sequences of human genes could be helpful. Distinguishing between real nucleotide sequences of a real gene and a mutated version seemed like an exciting task, since the ability to do such a task well requires some level of understanding of the data we were working on. In fact, what we did is similar to the evaluation of GPT-2 [36] on tasks like Winograd Schema Challenge [21], in the sense that our model is trained as a generative model, but it could be adopted to do discriminative tasks when the correct setting is prepared.
@@ 4.4.2 Real Mutation Dataset
Human Gene Mutation Database [41, 43, 42] was a nice resource with almost no alternative for our goal. Even though the number of mutations we could find for genes in our test set was only 10, which is very limited, it still helped to show that our models could make nice predictions for real examples as well. Since the database is not publicly available, we will, unfortunately, not be able to share the obtained mutation dataset.
@@ 6 Future Work
Acknowledgments
This work is partially supported by Boğaziçi University Research Fund under the Grant Number 16903 and by ERC grant (LifeLU, 101089287). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.